A car and van are parked 120 mi apart on a straight road. The drivers start driving toward each other at noon, each at the speed of 40 mi/h. A fly starts from the front bumper if the van at noon and flies to the bumper of the car, then immediately back to the bumper of the van, back to the car, and so on until the car and the van meet. If the fly flies at 100 mi/h, what is the total distance it travels?
Answer by jorel1380(3719)

The car and the van are 120 miles apart. They are approaching each other at 40 mph. Therefore, it takes them 120/40+40=120/80=3/2 hours to meet. The fly is flying at 100 mph for 3/2 hours, and thus therefore travels 100x3/2=150 miles between the two vehicles..before being squashed..